Summer Tyme, Inc., is considering a new 3-year expansion project that requires an initial fixed asset investment of $4.698 million. The fixed asset will be depreciated straight-line to zero over its 3-year tax life, after which time it will be worthless. The project is estimated to generate $4,176,000 in annual sales, with costs of $1,670,400. Required: If the tax rate is 35 percent, what is the OCF for this project? rev: 09_18_2012 $2,176,740 $610,740 $2,505,600 $2,067,903 $2,285,577

Dog Up! Franks is looking at a new sausage system with an installed cost of $241,800. This cost will be depreciated straight-line to zero over the project's 8-year life, at the end of which the sausage system can be scrapped for $37,200. The sausage system will save the firm $74,400 per year in pretax operating costs, and the system requires an initial investment in net working capital of $17,360.

  

Required:
If the tax rate is 33 percent and the discount rate is 15 percent, what is the NPV of this project?


rev: 09_18_2012

$26,641.67

$14,956.68

$23,104.38

$34,789.37

$24,259.60
Answer #1

1)

Annual depreciation = 4,698,000 / 3 = 1,566,000

Operating cash flow = ( sales - costs - depreciation)( 1 - tax) + depreciation

Operating cash flow = ( 4,698,000 - 1,670,400 - 1,566,000)( 1 - 0.35) + 1,566,000

Operating cash flow = $2,176,740

2)

Annual depreciation = 241,800 / 8 = 30,225

Initial investment = 241,800 + 17,360 = 259,160

Annual operating cash flow = (cash inflow - depreciation) ( 1 - tax) + depreciation

Annual operating cash flow = (74,400 - 30,225) ( 1 - 0.33) + 30,225

Annual operating cash flow = $59,822.25

Year 8 non-operating cash flow = Sale value + NWC - 0.33( sale - book value)

Year 8 non-operating cash flow = 37,200 + 17,360 - 0.33( 37,200 - 0)

Year 8 non-operating cash flow = 54,560 - 12,276

Year 8 non-operating cash flow = 42,284

Total year 8 cash inflow = 42,284 + 59,822.25 = 102,106.25

NPV = present value of cash onflows - present value of cash outflows

NPV = -259,160 + 59,822.25 / ( 1 + 0.15)1 + 59,822.25 / ( 1 + 0.15)2 + 59,822.25 / ( 1 + 0.15)3 + 59,822.25 / ( 1 + 0.15)4 + 59,822.25 / ( 1 + 0.15)5 + 59,822.25 / ( 1 + 0.15)6 + 59,822.25 / ( 1 + 0.15)7 + 102,106.25 / ( 1 + 0.15)8

NPV = $23,104.38